OpenStudy (anonymous):

slope intercept form the equation of the line passing through the points (2.3) (-4.5)

OpenStudy (campbell_st):

can you find the slope of the line..?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok... so what did you get for the slope...?

OpenStudy (anonymous):



OpenStudy (campbell_st):

ok... so the slope intercept form of a line is y = mx + b m is the slope and b is the y- intercept simplifying the slope it becomes - 1/3 so you have y = -1/3 x + b pick either point.... and substitute it into the equation to find b. if you choose (2, 3) substitute x = 2 and y = 3 into the equation and solve for b so 3 = -1/3 (2) + b can you do that..?

OpenStudy (anonymous):

ill try

OpenStudy (campbell_st):

try and leave your answer as a fraction

OpenStudy (campbell_st):

oops... it will be a mixed number... for b

OpenStudy (anonymous):